A solution or hydrogen peroxide is 26.3 H2O2 by mass and has a density of 1.11 g/cm3 the molarity of the solution is
Answer #1

answer :

molarity is defined as ratio of mole of H2O2 and volume of solution in liter.

molarity=\frac{mole of H2O2}{volume of solution(lit.)}

we can also write it as;

molarity=\frac{mass of H2O2(gm)}{m.w.of H2O2(gm)*volume of solution(lit.)}

now we multiply ans devide the formula by MASS OF SOLUTION IN GRAM;

molarity=\frac{mass of H2O2(gm)}{m.w.of H2O2(gm)*volume of solution(lit.)}*\frac{mass of solution in (gm)}{mass of solution in (gm)}

rearrange this equation;

molarity=\frac{mass of H2O2(gm)}{mass of solution in (gm)}*\frac{mass of solution in (gm)}{volume of solution(lit.)}*\frac{1}{m.w.of H2O2}

WE CAN ALSO WRITE IT AS;

molarity=percentageW/W*densityof solution(gm/lit)*\frac{1}{m.w.of H2O2}

given data :

mass percent of H2O2 = 26.3%

so %W/W = 26.3/100 = 0.263

density of solution = 1.11 gm/cm3=1.11 gm/ml = 1110 gm/lit

m.w of H2O2 = 34 gm/mole

molarity=percentageW/W*densityof solution(gm/lit)*\frac{1}{m.w.of H2O2}

molarity=0.263*1110*(gm/lit)*\frac{1}{34gm/mol}

molarity=8.59 mole/lit

molarity=8.59 molar